OpenStudy (anonymous):

transformation: x=2u y = 3v double integral (x^2 y) dA equation: 9x^2 + 4y^2

OpenStudy (watchmath):

Apa kabarnya IPB? :D

OpenStudy (anonymous):

9x^2 + 4y^2=36 36u^2 + 36v^2 = 36 36 (u^2+v^2) = 36 jacobian = 6 \[\int\limits_{o}^{2\pi}\int\limits_{0}^{1} 36r^2 r dr d \Theta\] is it true my double integral?

OpenStudy (anonymous):

woah ada orang indo, seperti biasa aja ipb, lagi rame penerimaan 48

OpenStudy (watchmath):

In the original integral what is A?

OpenStudy (anonymous):

area?

OpenStudy (watchmath):

I mean usually we compute the double integral over some area A (or S usually). What is the area? is it a cicle, area bounded by some curves or what?